two boy on bicycles start from he same place at the same time.one cyclist travel 15 miles per hour, the other travel 12 miles per hour. if they travel in the same direction how many hours will it take before they are 9 miles apart


Answer by ankor@dixie-net.com(22740)

two boy on bicycles start from he same place at the same time.
one cyclist travel 15 miles per hour, the other travel 12 miles per hour.
if they travel in the same direction how many hours will it take before
they are 9 miles apart
:
Let t = time for them to be 9 mi apart
:
Write a dist equation; dist = speed*time
:
15t - 12t = 9
3t = 9
t = 3 hrs
:
:
Confirm this, find the dist each travels:
15*3 = 45
12*3 = 36
----------
diff: 9 mi